OpenStudy (anonymous):

. What is 3.2 × 10−4 m/h in centimeters per second? A. 8.889 × 10−6 cm/s B. 8.889 × 10−5 cm/s C. 1.152 × 10−5 cm/s D. 1.152 × 10−7 cm/s

OpenStudy (anonymous):

Convert meters to centimeters

OpenStudy (math&ing001):

\[3.2*10^{-4} \frac{ meters }{ hour } = 3.2*10^{-4}\frac{ 100 cm }{ 3600 seconds }=3.2*10^{-4}*\frac{ 100 }{ 3600 } cm/s\]